OpenStudy (anonymous):

Help please he radius r of a circle is increasing at a rate of 7 centimeters per minute. (a) Find the rate of change of the area when r = 10 centimeters. (b) Find the rate of change of the area when r = 30 centimeters.

OpenStudy (anonymous):

A = pi r^2 dA/dt = 2pi r dr/dt plug and chug

OpenStudy (anonymous):

2*13.4*100=2680 right

ganeshie8 (ganeshie8):

you're given dr/dt = 7, dA/dt = 2 pi*r * 7 = 14pi*r

ganeshie8 (ganeshie8):

(a) Find the rate of change of the area when r = 10 centimeters. plugin r = 10 dA/dt = 14pi*10 = 140pi

ganeshie8 (ganeshie8):

similarly, see if u can do the part b)

OpenStudy (anonymous):

14pi*30=420pi right

ganeshie8 (ganeshie8):

Correct !

OpenStudy (anonymous):

thank youuuu